# Banach algebras associated to twisted étale groupoids: simplicity and pure infiniteness

### v2

@@ 6. Pure infiniteness
    for every $y\in B\setminus\{0\}$ there are $x,z\in B$ such that $xyz$ is an infinite idempotent in $B$.
@@
Condition ($\dagger$) holds if and only if every non-zero left (not necessarily closed) ideal in $B$ contains an infinite idempotent if and only if every non-zero right ideal in $B$ contains an infinite idempotent, see [AAM17, Proposition 3.1.7]. Thus ($\dagger$) is equivalent to apparently stronger and unsymmetrical conditions that for every $y\in B\setminus\{0\}$ there is $x\in B$ (resp. $z\in B$) such that $xy$ (resp. $yz$) is an infinite idempotent in $B$.
@@
      $B$ is a purely infinite in the sense of [Phi13a], [DH22], i.e. $B\not\cong\mathbb{F}$ and for every $y\in B\setminus\{0\}$ there are $x,z\in B$ such that $xyz=1$.
@@
      $B$ is a purely infinite in the sense of [CMR23], i.e. $B\not\cong\mathbb{F}$ and for every $a,b\in B\setminus\{0\}$ there are sequences $\{x_{n}\}_{n=1}^{\infty},\{z_{n}\}_{n=1}^{\infty}$ in $B$ such that $x_{n}bz_{n}\to a$;
@@
      for every idempotent $e\in B\setminus\{0\}$ the unital Banach algebra $eBe$ is purely infinite in the sense of the equivalent conditions in (1).
@@
    If $B$ is a $C^{*}$-algebra then $B$ is a purely infinite Banach algebra if and only if it is purely infinite in the sense of Cuntz [Cun81], i.e. every non-zero hereditary $C^{*}$-subalgebra of $B$ contains a partial isometry $u$ with $uu^{*}<u^{*}u$.
@@
(1). Conditions (1)(a) and (1)(b) differ only in that one assumes algebraic simplicity and the other analytic simplicity, but for unital Banach algebras these are equivalent by openness of the set of invertible elements. The equivalence (1)(b)$\Leftrightarrow$(1)(c) holds by [AGP02, Theorem 1.6].
(2). Implications (2)(a)$\Rightarrow$(2)(c) and (2)(b)$\Rightarrow$(2)(c) hold in general. Indeed, for (2)(b)$\Rightarrow$(2)(c) see [CMR23, Corollary 3.4], and for (2)(a)$\Rightarrow$(2)(c) assume (2)(a) and let $e\in B\setminus\{0\}$ be an idempotent. Then $eBe$ is a unital Banach algebra, and for any nonzero $y\in eBe$ there is $x\in B$ such that $xy\in Be$ is an infinite projection in $Be$. Then $xy=xye\sim exy=exey\in eBe$, and so for $x^{\prime}:=exe\in eBe$ the product $x^{\prime}y$ is an infinite projection in $eBe$. Hence $eBe$ purely infinite, see Remark 6.2.
Now let us assume (2)(c) and that $B$ has a bounded two-sided approximate unit consisting of idempotents. Then $B\neq\mathbb{F}$. Let $a,b\in B\setminus\{0\}$. Then we may choose a sequence $\{e_{n}\}$ of idempotents in $B$ such that $e_{n}ae_{n}\to a$ and $e_{n}be_{n}\to b$, and by (2)(c) for each $n$ we may find $x_{n},z_{n}\in e_{n}Be_{n}$ such that $x_{n}(e_{n}be_{n})y_{n}=e_{n}ae_{n}$. Then $x_{n}bz_{n}\to a$, which proves (2)(b). Now let $y\in B\setminus\{0\}$. Then there is an idempotent $e\in B$ such that $eye\in eBe$ is nonzero. By (2)(c) there are $x,z\in eBe$ such that $xeyez$ is an infinite idempotent in $eBe$. This proves (2)(a).
(3). Assume $B$ is purely infinite in the sense of Cuntz [Cun81]. Let $y\in B\setminus\{0\}$. We may assume that $\|y\|=1$. Put $a:=y^{*}y$ and let $f$ be a continuous function which takes the value 00 on $(-\infty,1/2]$ and the value $1$ on $[1,+\infty)$, and is linear on $[1/2,1]$. Then $af(a)\geq f(a)/2>0$ in $\overline{aBa}\subseteq\overline{y^{*}By}$. By assumption there is a partial isometry $u\in\overline{f(a)Bf(a)}$ such that $uu^{*}<u^{*}u$. Writing $u=f(a)b$ with $b\in\overline{f(a)Bf(a)}$ we get $u^{*}au=b^{*}f(a)af(a)b\geq 1/2\cdot b^{*}f(a)f(a)b=1/2\cdot u^{*}u$. Thus $u^{*}au$ is invertible in the unital algebra $\overline{u^{*}Bu}$ with $u^{*}u$ being a unit. Denote the inverse of $u^{*}au$ by $z$, and putting $w:=zu^{*}a=zu^{*}y^{*}y\in By$. Then $uw\in By$ is an idempotent equivalent to the infinite idempotent $wu=u^{*}u$. Hence $uw\in By$ is infinite.
Conversely, assume $B$ is purely infinite in the sense of Definition 6.1 and let $y\in B\setminus\{0\}$. Then there is an infinite idempotent $e$ in $By$. Every idempotent $e$ in a $C^{*}$-algebra is similar (in fact even homotopy equivalent) to a projection $p$ with $p=pe$, see the proof of [Bla98, Proposition 4.6.2]. Since $p\in Be\subseteq By$ and $p$ is self-adjoint we get $p\in\overline{y^{*}By}$. Since $e$ is an infinite element, it follows that $p$ is an infinite projection. Indeed, $p$ is an infinite element because it is equivalent to an infinite element $e$, and so there is an idempotent $f$ with $p\sim f<p$. Then, as mentioned above, there is a projection $q\sim f$ with $q=qf$. Thus we have $q\sim p$ in $\overline{y^{*}By}$ and so they are Murray–von Neumann equivalent, see [Bla98, Proposition 4.6.4]. Using that $f<p$ and $q=qf$ we get $q=qf=q(fp)=(qf)p=qp$. Therefore $q\leq p$ (because $p$ and $q$ are self-adjoint) and in fact $q<p$ (because $q=p$ would imply $f=p$). ∎
@@
Without the standing assumption in (2) above, we do not know the relationship between pure infiniteness in our sense and that of [CMR23]. When $B$ is a simple unital $C^{*}$-algebra the equivalence between the Cuntz definition and the condition in (1)(b) is well known, see [Bla98, Proposition 6.11.5].
@@
Let $\mathcal{G}$ be $n$-filling with respect to a unital inverse semigroup $S\subseteq S(\mathcal{L})$ covering $\mathcal{G}$. For any $b\in C(X)^{+}$ and any $\varepsilon>0$ there is $c\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ such that $\|c\|_{\max}^{S}\leq n$, $c^{*}*b*c\in C(X)$ and $c^{*}*b*c\geq\|b\|_{\infty}-\varepsilon$.
@@
We may assume that $\|b\|_{\infty}=1$. Using the $n$-filling property, exactly as in the proof of [Suz17, Lemma 3.10] one may construct mutually disjoint non-empty subsets $U_{1},\ldots,U_{n}$ of $U:=\{x\in X:b(x)\geq 1-\varepsilon\}$, relatively compact bisections $W_{1},\dots W_{n}\in\operatorname{Bis}(\mathcal{G})$, where each $W_{i}$ is a subset of an element of $S$, and closed subsets $K_{i}\subseteq W_{i}$, $i=1,\ldots,n$, such that
@@
Now, since the bundle over each $W_{i}$ can be trivialized, we may pick sections $f_{i}\in C_{c}(W_{i},\mathcal{L})$ such that $\|f_{i}\|_{\infty}=1$ and $|f_{i}|_{K_{i}}|=1$, for $i=1,\ldots,n$. Putting $c:=\sum_{i=1}^{n}f_{i}^{*}$ we have $\|c\|_{\max}^{S}\leq\sum_{i=1}^{n}\|f_{i}^{*}\|_{\infty}=n$, and since $\{W_{i}\}_{i=1}^{n}$ are bisections with mutually disjoint domains $\{d(W_{i})\}_{i=1,\ldots,n}$ we get $c^{*}*b*c=\sum_{i=1}f_{i}^{*}*b*f_{i}\in C(X)$. Since $d(W_{i})\subseteq U$, for $i=1,\ldots,n$ and $\bigcup_{i=1}^{n}r(K_{i})=X$, we further obtain $c^{*}*b*c\geq 1-\varepsilon$. ∎
@@
Assume that $\mathcal{G}$ is topologically free and $n$-filling with respect to a unital inverse semigroup $S\subseteq S(\mathcal{L})$ covering $\mathcal{G}$. Every $S$-graded essential Banach algebra $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ is purely infinite simple (and unital).
@@
Let $y\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\setminus\{0\}$. By Proposition 6.3(1) it suffices to show that there are $x,z\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ such that $xyz$ is invertible. By Proposition 4.12, there is $U\in S(\mathcal{L})$ and $a_{U}\in C_{c}(r(U))$ such that $\operatorname{\mathbb{E}}^{\mathcal{R}}_{U}(a_{U}y)\neq 0$. Replacing $y$ with $a_{U}y$ and scaling, we may assume that $y\in C_{c}(r(U))F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ and $\|\operatorname{\mathbb{E}}^{\mathcal{R}}_{U}(y)\|=1$. Let $\varepsilon>0$ be such that $\varepsilon<1/(2n^{2}+2)$. Let $i_{\mathcal{R}}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ be the canonical homomorphism, which is isometric on the subspaces $C_{c}(V,\mathcal{L})$, $V\in S$. The argument from the beginning of the proof of [JR00, Theorem 1.2] shows that we may find $f\in C_{c}(r(U))\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ with $\|y-i_{\mathcal{R}}(f)\|_{\mathcal{R}}<\varepsilon$ and $\|\operatorname{\mathbb{E}}^{\mathcal{R}}_{U}(i_{\mathcal{R}}(f))\|=\|q_{U}(f\circ r|_{U}^{-1})\|=1$. By Lemma 3.5 there is a contractive $b\in C_{c}(U^{*},\mathcal{L})$ such that $|b*f|_{X}|=|f\circ r|_{U}^{-1}|$. Thus replacing $y$ with $by$ and $f$ with $bf$ we may assume that $U=X$, i.e. we have $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ with
@@
In particular, $\|a(bf|_{X})a\|\geq 1-\varepsilon$. Applying Lemma 6.6 to $a(bf|_{X})a$ we get $c\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ such that $\|c\|_{\max}^{S}\leq n$ and $g:=c^{*}(a(bf|_{X})a)c\in C(X)$ satisfies $g\geq 1-2\varepsilon$. In particular, $d:=i_{\mathcal{R}}(c)\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ satisfies $\|d\|_{\mathcal{R}}\leq\|c\|_{\max}^{S}\leq n$ and $g$ is invertible in $C(X)\subseteq F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ where $\|g^{-1}\|_{\infty}\leq(1-2\varepsilon)^{-1}$. Thus every $h\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ with $\|h-g\|_{\mathcal{R}}<1-2\varepsilon$ is invertible in $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$. Using the above displayed inequalities, together with $\|d\|_{\mathcal{R}}\leq n$, $\|a\|,\|b\|\leq 1$, and our choice of $\varepsilon$ we get
@@
If $S=S(\mathcal{L})=\operatorname{Bis}(\mathcal{G})$, so for instance when the twist comes from a $2$-cocycle (equivalently $\mathcal{L}$ is trivial as a line bundle, see Example 2.2), then the properties in Definitions 6.56.8 do not depend on the twist, and if they hold we simply call the groupoid $\mathcal{G}$ $n$-filling [Suz17] or locally contractive [A-D97], respectively.
@@
The $n$-filling property implies that $\mathcal{G}$ is minimal, while locally contractive does not. Assuming minimality, the relationship between $n$-filling and locally contracting is not completely clear. However, when $\mathcal{G}$ is ample it is known that $n$-filling does not depend on $n$ (is equivalent to $1$-filling) and it implies locally contracting, see [Ma22, Theorem 5.8] (it also implies paradoxicality of all non-empty compact open sets, see [RS20, Proposition 7.7]).
@@
Assume $\mathcal{G}$ is locally contracting with respect to a unital inverse semigroup $S\subseteq S(\mathcal{L})$ covering $\mathcal{G}$. For any bump function $a\in C_{c}(X)$ there is an idempotent $p\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ with $p=ap=pa$ and $\|p\|_{\max}^{S}\leq 4$, and there is $u\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, with $\|u\|_{\max}^{S}\leq 8$, such that $u^{*}u=p$ and $uu^{*}<p$ (so $p$ is infinite). Moreover, these relations remain valid for the images of $u$ and $p$ in any $S$-graded essential Banach algebra $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ of $(\mathcal{G},\mathcal{L})$.
@@
We adapt the argument from the proof of [A-D97, Proposition 2.2], see also [Sim20, Theorem 4.4.2]. If $X$ is not compact we let $\widetilde{\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})}$ be the minimal unitization of the algebra $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$. Then $1\mathbb{C}+C_{c}(X)=\widetilde{C_{c}(X)}\subseteq\widetilde{\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})}$ is naturally a subalgebra and $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\subseteq\widetilde{\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})}$ is a two-sided ideal. Moreover, every positive function in $\widetilde{C_{c}(X)}$ has a square root in $\widetilde{C_{c}(X)}$.
Let $a\in C_{c}(X)$ be a bump function. Let $U$ be a non empty open set such that $a|_{U}\equiv 1$. Take $V\subseteq U$ and $W\in S$ with $\overline{V}\subseteq d(W)$ and $h_{W}(\overline{V})\subsetneq V$. Find an open $V_{0}$ such that $h_{W}(\overline{V})\subseteq V_{0}\subseteq\overline{V_{0}}\subsetneq V$. Fix a positive norm one function $b\in C_{c}(V_{0})$ which is equal to $1$ on $h_{W}(\overline{V})$. Then $1-b^{2}\in\widetilde{C_{c}(X)}$ vanishes on $h_{W}(\overline{V})$ and $b\circ h_{W^{*}}\in C_{c}(h_{W}(V_{0}))$, so in particular $(1-b^{2})^{1/2}(b\circ h_{W^{*}})=0$. Since the bundle $\mathcal{L}$ can be trivialized on $W$ we may pick a unitary section $c_{W}\in C(W,\mathcal{L})$. Put $x:=c_{W}b=(b\circ h_{W^{*}})c_{W}$. Then $x^{*}x=b^{2}\in C_{c}(V)$ and $xx^{*}=b^{2}\circ h_{W^{*}}\in C_{c}(h_{W}(V))$. It follows that $(1-x^{*}x)^{1/2}x=0$ ($x$ is the so called scaling element). Therefore
@@
To prove that in fact $i_{\mathcal{R}}(p)$ is infinite we take any bump function $c\in C_{c}(X)$ with support in $O$, so that we get $cx^{*}x=c$ and $cx=0$. Applying the above reasoning to $c$ (instead of $a$) we get a scaling element $y$ with $cy=yc=y$. This implies $yx^{*}x=x^{*}xy=y$ and $yx=0$. Hence $py=yp=y$. As a consequence $p$ commutes with the elements in the Banach subalgebra of $\widetilde{C_{c}(X)}$ generated by the positive function $1-y^{*}y$, and in particular with $(1-y^{*}y)^{1/2}$. Thus $p$ commutes with $w:=y+(1-y^{*}y)^{1/2}\in\widetilde{\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})}$. Therefore putting $u:=pw\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, and using that $w^{*}w=1$ (which holds because $y$ is scaling), we get $u^{*}u=p$ and
@@
Assume $\mathcal{G}$ is topologically free and locally contracting with respect to a unital inverse semigroup $S\subseteq S(\mathcal{L})$ covering $\mathcal{G}$. Every non-zero left ideal in every $S$-graded essential Banach algebra $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ contains an infinite idempotent.

### v3

@@ 6. Pure infiniteness
    for every $y\in B\setminus\{0\}$, there are $x,z\in B$ such that xyzxyz is an infinite idempotent in $B$.
@@
Condition ($\dagger$) holds if and only if every non-zero left ideal (not necessarily closed) in $B$ contains an infinite idempotent, see [AAM17, Proposition 3.1.7]. Thus ($\dagger$) is equivalent to the apparently stronger and unsymmetrical condition that, for every $y\in B\setminus\{0\}$ there is $x\in B$ such that $xy$ is an infinite idempotent in $B$. A version of this remark where left ideals are replaced by right ideals is also true.
@@
      $B$ is purely infinite in the sense of [Phi13a], [DH22], that is, $B\not\cong\mathbb{F}$ and for every $y\in B\setminus\{0\}$ there are $x,z\in B$ such that $xyz=1$.
@@
      $B$ is purely infinite in the sense of [CMR23], that is, $B\not\cong\mathbb{F}$ and for every $a,b\in B\setminus\{0\}$ there are sequences $(x_{n})_{n=1}^{\infty},(z_{n})_{n=1}^{\infty}$ in $B$ such that $x_{n}bz_{n}\to a$;
@@
      for every idempotent $e\in B\setminus\{0\}$, the unital Banach algebra $eBe$ is purely infinite in the sense of the equivalent conditions in (1).
@@
    If $B$ is a $C^{*}$-algebra, then $B$ is a purely infinite Banach algebra if and only if it is purely infinite in the sense of Cuntz [Cun81], that is, every non-zero hereditary $C^{*}$-subalgebra of $B$ contains a partial isometry $u$ with $uu^{*}<u^{*}u$.
@@
[…]
Now let us assume (2)(c) and that $B$ has a bounded two-sided approximate unit consisting of idempotents. By assumption $B\neq\mathbb{F}$. Let $a,b\in B\setminus\{0\}$. Then we may choose a sequence $(e_{n})_{n}$ of idempotents in $B$ such that $e_{n}ae_{n}\to a$ and $e_{n}be_{n}\to b$, and by (2)(c) for each $n$ we may find $x_{n},z_{n}\in e_{n}Be_{n}$ such that $x_{n}(e_{n}be_{n})y_{n}=e_{n}ae_{n}$. Then $x_{n}bz_{n}\to a$, which proves (2)(b). Now let $y\in B\setminus\{0\}$. Then there is an idempotent $e\in B$ such that $eye\in eBe$ is nonzero. By (2)(c) there are $x,z\in eBe$ such that xeyezxeyez is an infinite idempotent in $eBe$. This proves (2)(a).
(3). Assume $B$ is purely infinite in the sense of Cuntz [Cun81]. Let $y\in B\setminus\{0\}$. We may assume that $\|y\|=1$. Put $a:=y^{*}y$ and let $f$ be a continuous function which takes the value 0 on $(-\infty,1/2]$ and the value $1$ on $1,+\infty)$, and is linear on $[1/2,1]$. Then $af(a)\geq f(a)/2>0$ in $\overline{aBa}\subseteq\overline{y^{*}By}$. By assumption, there is a partial isometry $u\in\overline{f(a)Bf(a)}$ such that $uu^{*}<u^{*}u$. Writing $u=f(a)b$ with $b\in\overline{f(a)Bf(a)}$, we get $u^{*}au=b^{*}f(a)af(a)b\geq 1/2\cdot b^{*}f(a)f(a)b=1/2\cdot u^{*}u$. Thus $u^{*}au$ is invertible in the unital algebra $\overline{u^{*}Bu}$ with $u^{*}u$ being a unit. Denote the inverse of $u^{*}au$ by $z$, and put $w:=zu^{*}a=zu^{*}y^{*}y\in By$. Then $uw\in By$ is an idempotent equivalent to the infinite idempotent $wu=u^{*}u$. Hence $uw\in By$ is infinite.
Conversely, assume $B$ is purely infinite in the sense of Definition 6.1 and let $y\in B\setminus\{0\}$. Then there is an infinite idempotent $e$ in $By$. Every idempotent $e$ in a $C^{*}$-algebra is similar (in fact even homotopy equivalent) to a projection $p$ with $p=pe$, see the proof of [Bla98, Proposition 4.6.2]. Since $p\in Be\subseteq By$ and $p$ is self-adjoint, we get $p\in\overline{y^{*}By}$. Since $e$ is an infinite element, it follows that $p$ is an infinite projection. Indeed, $p$ is an infinite element because it is equivalent to an infinite element $e$, and so there is an idempotent $f$ with $p\sim f<p$. Then, as mentioned above, there is a projection $q\sim f$ with $q=qf$. Thus we have $q\sim p$ in $\overline{y^{*}By}$ and so they are Murray–von Neumann equivalent, see [Bla98, Proposition 4.6.4]. Using that $f<p$ and $q=qf$, we get $q=qf=q(fp)=(qf)p=qp$. Therefore $q\leq p$ (because $p$ and $q$ are self-adjoint) and in fact $q<p$ (because $q=p$ would imply $f=p$). ∎
@@
Without the standing assumption in (2) above, we do not know the relationship between pure infiniteness in our sense and that of [CMR23]. When $B$ is a simple unital $C^{*}$-algebra, the equivalence between the Cuntz definition and the condition in (1)(b) is well known, see [Bla98, Proposition 6.11.5].
@@
Let $\mathcal{G}$ be $n$-filling with respect to a unital inverse semigroup $S\subseteq S(\mathcal{L})$ covering $\mathcal{G}$. For any $b\in C(X)^{+}$ and any $\varepsilon>0$, there is $c\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ such that $\|c\|_{\max}^{S}\leq n$, $c^{*}*b*c\in C(X)$ and $c^{*}*b*c\geq\|b\|_{\infty}-\varepsilon$.
@@
We may assume that $\|b\|_{\infty}=1$. Using the $n$-filling property, exactly as in the proof of [Suz17, Lemma 3.10], one may construct mutually disjoint non-empty subsets $U_{1},\ldots,U_{n}$ of $U:=\{x\in X:b(x)\geq 1-\varepsilon\}$, relatively compact bisections $W_{1},\ldots,W_{n}\in\operatorname{Bis}(\mathcal{G})$, where each $W_{i}$ is a subset of an element of $S$, and closed subsets $K_{i}\subseteq W_{i}$, $i=1,\ldots,n$, such that
@@
Now, since the bundle over each $W_{i}$ can be trivialized, we may pick sections $f_{i}\in C_{c}(W_{i},\mathcal{L})$ such that $\|f_{i}\|_{\infty}=1$ and $|f_{i}|_{K_{i}}|=1$, for $i=1,\ldots,n$. Putting $c:=\sum_{i=1}^{n}f_{i}^{*}$, we have $\|c\|_{\max}^{S}\leq\sum_{i=1}^{n}\|f_{i}^{*}\|_{\infty}=n$, and since $\{W_{i}\}_{i=1}^{n}$ are bisections with mutually disjoint domains $\{d(W_{i})\}_{i=1,\ldots,n}$, we get $c^{*}*b*c=\sum_{i=1}f_{i}^{*}*b*f_{i}\in C(X)$. Since $d(W_{i})\subseteq U$, for $i=1,\ldots,n$ and $\bigcup_{i=1}^{n}r(K_{i})=X$, we further obtain $c^{*}*b*c\geq 1-\varepsilon$. ∎
@@
Assume that $\mathcal{G}$ is topologically free and $n$-filling with respect to a unital inverse semigroup $S\subseteq S(\mathcal{L})$ covering $\mathcal{G}$. Then every $S$-graded essential Banach algebra $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ is purely infinite simple (and unital).
@@
Let $y\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\setminus\{0\}$. By Proposition 6.3(1), it suffices to show that there are $x,z\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ such that xyzxyz is invertible. By Proposition 4.12, there is $U\in S(\mathcal{L})$ and $a_{U}\in C_{c}(r(U))$ such that $\operatorname{\mathbb{E}}^{\mathcal{R}}_{U}(a_{U}y)\neq 0$. Replacing $y$ with $a_{U}y$ and scaling, we may assume that $y\in C_{c}(r(U))F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ and $\|\operatorname{\mathbb{E}}^{\mathcal{R}}_{U}(y)\|=1$. Let $\varepsilon>0$ be such that $\varepsilon<1/(2n^{2}+2)$. Let $i_{\mathcal{R}}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ be the canonical homomorphism, which is isometric on the subspaces $C_{c}(V,\mathcal{L})$, $V\in S$. The argument from the beginning of the proof of [JR00, Theorem 1.2] shows that we may find $f\in C_{c}(r(U))\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ with $\|y-i_{\mathcal{R}}(f)\|_{\mathcal{R}}<\varepsilon$ and $\|\operatorname{\mathbb{E}}^{\mathcal{R}}_{U}(i_{\mathcal{R}}(f))\|=\|q_{U}(f\circ r|_{U}^{-1})\|=1$. By Lemma 3.5, there is a contractive $b\in C_{c}(U^{*},\mathcal{L})$ such that $|b*f|_{X}|=|f\circ r|_{U}^{-1}|$. Thus, replacing $y$ with $by$ and $f$ with $bf$, we may assume that $U=X$, that is, we have $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ with
@@
In particular, $\|a(bf|_{X})a\|\geq 1-\varepsilon$. Applying Lemma 6.6 to $a(bf|_{X})a$, we get $c\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ such that $\|c\|_{\max}^{S}\leq n$ and $g:=c^{*}(a(bf|_{X})a)c\in C(X)$ satisfies $g\geq 1-2\varepsilon$. In particular, $d:=i_{\mathcal{R}}(c)\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ satisfies $\|d\|_{\mathcal{R}}\leq\|c\|_{\max}^{S}\leq n$ and $g$ is invertible in $C(X)\subseteq F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ where $\|g^{-1}\|_{\infty}\leq(1-2\varepsilon)^{-1}$. Thus every $h\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ with $\|h-g\|_{\mathcal{R}}<1-2\varepsilon$ is invertible in $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$. Using the above displayed inequalities, together with $\|d\|_{\mathcal{R}}\leq n$, $\|a\|,\|b\|\leq 1$, and our choice of $\varepsilon$, we get
@@
If $S=S(\mathcal{L})=\operatorname{Bis}(\mathcal{G})$, so for instance when the twist comes from a $2$-cocycle (equivalently $\mathcal{L}$ is trivial as a line bundle, see Example 2.2), then the properties in Definitions 6.5 and 6.8 do not depend on the twist, and if they hold, we simply call the groupoid $\mathcal{G}$ $n$-filling [Suz17] or locally contractive [A-D97], respectively.
@@
The $n$-filling property implies that $\mathcal{G}$ is minimal, while locally contracting does not. Assuming minimality, the relationship between $n$-filling and locally contracting is not completely clear. However, when $\mathcal{G}$ is ample, it is known that $n$-filling does not depend on $n$ (is equivalent to $1$-filling) and it implies locally contracting, see [Ma22, Theorem 5.8] (it also implies paradoxicality of all non-empty compact open sets, see [RS20, Proposition 7.7]).
@@
Assume that $\mathcal{G}$ is locally contracting with respect to a unital inverse semigroup $S\subseteq S(\mathcal{L})$ covering $\mathcal{G}$. For any bump function $a\in C_{c}(X)$, there is an idempotent $p\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ with $p=ap=pa$ and $\|p\|_{\max}^{S}\leq 4$, and there is $u\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, with $\|u\|_{\max}^{S}\leq 8$, such that $u^{*}u=p$ and $uu^{*}<p$ (so $p$ is infinite). Moreover, these relations remain valid for the images of $u$ and $p$ in any $S$-graded essential Banach algebra $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ of $(\mathcal{G},\mathcal{L})$.
@@
We adapt the argument from the proof of [A-D97, Proposition 2.2], see also [Sim20, Theorem 4.4.2]. If $X$ is not compact, we let $\widetilde{\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})}$ be the minimal unitization of the algebra $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$. Then $1\mathbb{C}+C_{c}(X)=\widetilde{C_{c}(X)}\subseteq\widetilde{\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})}$ is naturally a subalgebra and $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\subseteq\widetilde{\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})}$ is a two-sided ideal. Moreover, every positive function in $\widetilde{C_{c}(X)}$ has a square root in $\widetilde{C_{c}(X)}$.
Let $a\in C_{c}(X)$ be a bump function. Let $U$ be a non empty open set such that $a|_{U}\equiv 1$. Take $V\subseteq U$ and $W\in S$ with $\overline{V}\subseteq d(W)$ and $h_{W}(\overline{V})\subsetneq V$. Find an open $V_{0}$ such that $h_{W}(\overline{V})\subseteq V_{0}\subseteq\overline{V_{0}}\subsetneq V$. Fix a positive norm one function $b\in C_{c}(V_{0})$ which is equal to $1$ on $h_{W}(\overline{V})$. Then $1-b^{2}\in\widetilde{C_{c}(X)}$ vanishes on $h_{W}(\overline{V})$ and $b\circ h_{W^{*}}\in C_{c}(h_{W}(V_{0}))$, so in particular $(1-b^{2})^{1/2}(b\circ h_{W^{*}})=0$. Since the bundle $\mathcal{L}$ can be trivialized on $W$, we may pick a unitary section $c_{W}\in C(W,\mathcal{L})$. Put $x:=c_{W}b=(b\circ h_{W^{*}})c_{W}$. Then $x^{*}x=b^{2}\in C_{c}(V)$ and $xx^{*}=b^{2}\circ h_{W^{*}}\in C_{c}(h_{W}(V))$. It follows that $(1-x^{*}x)^{1/2}x=0$ ($x$ is the so called scaling element). Therefore
@@
To prove that in fact $i_{\mathcal{R}}(p)$ is infinite we take any bump function $c\in C_{c}(X)$ with support in $O$, so that we get $cx^{*}x=c$ and $cx=0$. Applying the above reasoning to $c$ (instead of $a$), we get a scaling element $y$ with $cy=yc=y$. This implies $yx^{*}x=x^{*}xy=y$ and $yx=0$. Hence $py=yp=y$. As a consequence $p$ commutes with the elements in the Banach subalgebra of $\widetilde{C_{c}(X)}$ generated by the positive function $1-y^{*}y$, and in particular with $(1-y^{*}y)^{1/2}$. Thus $p$ commutes with $w:=y+(1-y^{*}y)^{1/2}\in\widetilde{\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})}$. Therefore putting $u:=pw\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, and using that $w^{*}w=1$ (which holds because $y$ is scaling), we get $u^{*}u=p$ and
@@
Assume $\mathcal{G}$ is topologically free and locally contracting with respect to a unital inverse semigroup $S\subseteq S(\mathcal{L})$ covering $\mathcal{G}$. Then every non-zero left ideal in every $S$-graded essential Banach algebra $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ contains an infinite idempotent.